OpenStudy (anonymous):

Complete the Square: 4x^2 - 10x - 9 = 0 4(x^2 - 10x/4 - 9/4)=0 therefore if 4(0)=0 x^2 - 10x/4 - 9/4 = 0 (x - 10/4)^2 - 100/16 - 36/16 (x - 10/4)^2 - 136/16 Where do I go from here? Am I doing it right?

OpenStudy (anonymous):

no unfortunately

OpenStudy (anonymous):

If you could help me on the right track on getting X=A OR X=B

OpenStudy (anonymous):

\[x^2-\frac{5}{2}x=\frac{9}{4}\] \[(x-\frac{5}{4})^2=\frac{9}{4}+\frac{25}{16}\] etc

OpenStudy (anonymous):

first off i would write \[\frac{5}{2}\] instead of \[\frac{10}{4}\] but more important is that when you "complete the square" you have to take HALF the middle term to put inside the parentheses